# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v2

@@ 4.2 Edge selection and sign consistency
With further assumptions, the elementwise ($\ell_{\infty}$) error bound of $\hat{\boldsymbol{\Omega}}$ can be controlled with which edge selection and sign consistency can be proved. The Hessian matrix of the population risk $R(\boldsymbol{\Omega})$ (see §3) with respect to the usual vectorization of $\boldsymbol{\Omega}$ at $\boldsymbol{\Omega}^{*}$ is
$\boldsymbol{\Gamma}^{\ast}=\boldsymbol{\Sigma}^{*}\otimes\boldsymbol{I}_{p}+(\boldsymbol{\Omega}_{D}^{\ast-1}\otimes\boldsymbol{\Omega}_{D}^{\ast-1})\boldsymbol{\Upsilon},$ (4.1)
where $\otimes$ is a Kronecker product and $\boldsymbol{\Upsilon}=\sum_{i=1}^{p}e_{i}e_{i}^{T}\otimes e_{i}e_{i}^{T}$. It is positive definite since so is $\boldsymbol{\Sigma}^{\ast}$. Let $S$ be the support of $\boldsymbol{\Theta}^{*}$ as stated in Theorem 4.1.
Assumption 4.1.
We further assume the following properties on the data distribution.
  1. (A1)
    The maximum degree $d=\max_{i\in[p]}\left|\{j\in[p]:\theta^{\ast}_{ij}\neq 0\}\right|=\max_{i\in[p]}\left|\{j\in[p]|\omega^{\ast}_{ij}\neq 0\}\right|$ of the graph implied by $\boldsymbol{\Theta}^{*}$ is bounded.
  2. (A2)
    (Irrepresentability condition) For some $\alpha\in[0,1)$, the matrix $\boldsymbol{\Gamma}^{*}$ satisfies
@@
Let us define the following associated quantities: $\kappa_{\Gamma^{\ast}}=\vvvert(\boldsymbol{\Gamma}^{\ast}_{SS})^{-1}\vvvert_{\infty}$, $\kappa_{\Sigma^{\ast}}=\vvvert\boldsymbol{\Sigma}^{\ast}\vvvert_{\infty}$, $\gamma_{1}=\vvvert\boldsymbol{\Omega}^{\ast-1}_{D}\vvvert_{\infty}$. Recall $\kappa_{\Omega^{\ast}}=\vvvert\boldsymbol{\Omega}^{\ast}\vvvert_{\infty}$.
@@
    there holds $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{\ast}\|_{\infty}\leq 24\sqrt{2}(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\kappa_{\Gamma^{\ast}}\kappa_{\Omega^{\ast}}(1+10/\alpha)\sqrt{(\tau\log{p}+\log{4})/n}$;
@@
With other quantities held fixed, the sample size required to achieve the rate $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{*}\|_{\infty}=O_{P}(\sqrt{(\tau\log p)/n})$ is $n\gtrsim d^{2}\tau\log p$. It can be also shown that for data distributions with a bounded $4m$-th moment, the error rate is $O_{P}(\sqrt{p^{\tau/m}/n})$ for $n\gtrsim d^{2}p^{\tau/m}$; see Supplementary Material A.6. These sample complexities and rates of convergence match those of graphical lasso, obtained by Ravikumar et al. [33] under similar conditions.
@@
where $\sigma$ and $\delta$ are as defined in Theorem 4.2, then the event $\{\operatorname{sign}(\omega_{ij}^{*})=\operatorname{sign}(\hat{\omega}_{ij})$ for all $(i,j)\in S\}$ occurs with a probability no smaller than $1-p^{-(\tau-2)}$.
@@ 5.1 Linear convergence of ACCORD-FBS
ACCORD-FBS was investigated with two variants: one with a constant step size of $\tau={1}/{L}$ and the other employing backtracking line search with $\tau={1}/{L}$ as a lower bound. In contrast, a grid of constant step sizes, $\tau\in\{0.25,0.6,0.75\}$, along with backtracking, were chosen for ACCORD-ISTA to demonstrate its convergence behavior.
@@ 5.3 Edge detection and estimation error
We examine the edge detection and parameter estimation performance of ACCORD on simulated data, comparing it with CONCORD, graphical lasso (GLASSO), SPACE, CLIME. Following the hub graph setting in Peng et al. [32], we generated a graph with 10 clusters each having 100 nodes and 90 edges and connected 100 randomly chosen node pairs from adjacent clusters. The clusters had either a hub network or scale-free graph structure. A hub network was generated by (1) constructing an Erdös-Rényi graph with 97 nodes and 45 edges, and (2) consecutively adding 3 hub nodes, each connected to random 15 nodes. A scale-free graph was generated so that the degree $k$ of the nodes has a probability $P(k)\sim k^{-2.3}$. We additionally considered a simple Erdös-Rényi graph with 1,000 nodes and 1,000 edges without the clustered structure. The partial correlation matrix corresponding to the graph was chosen so that the matrix has minimum eigenvalue of at least 0.2, and each nonzero entries have absolute value of 0.1 at least. Model selection was carried out using epBIC (3.7) for ACCORD and CONCORD for an evenly spaced grid of $\lambda$’s in logarithmic scale, while extended Gaussian BIC was used for GLASSO and CLIME. Although cross validation was used to select the $\lambda$ in the original paper of CLIME [4], we used BIC instead to avoid the repetitive parameter fitting. For SPACE, we used “BIC-type criterion” defined in Peng et al. [32]. For CONCORD, the matrix $\hat{\boldsymbol{\Theta}}$ minimizing (2.2) was treated as a precision matrix despite the possible inconsistency (see §2). This procedure was repeated 50 times. For ACCORD, we also considered debiasing procedure (3.8).
@@
Table 1 reports the area under the precision-recall curve (AUPRC) and number of edges selected. In terms of AUPRC (and Matthews correlation coefficient; see Supplementary Material B), ACCORD performed slightly better than or similarly to other methods. The trend of the precision-recall curves did not particularly vary among replications. In all cases, the penalty selected from each selection method yielded reasonable number of selected edges. We could observe that BIC-type methods with the extended term drastically reduces false positive (FP) edges in the expense of few true positive (TP) edges.
Table 2 reports the total squared error of the estimated precision matrix $\boldsymbol{\Theta}$ and its reparameterization $\boldsymbol{\Omega}$ by ACCORD and CONCORD, along with their debiased refit. Compared to CONCORD, ACCORD estimates clearly showed better result for both $\hat{\boldsymbol{\Theta}}$ and $\hat{\boldsymbol{\Omega}}$ in terms of the estimate error. The debiased refit also clearly improves the estimating performance. For more details about the impact of the debiased refit, see Supplementary Material B.
@@ 6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
In this section, we show that HP-ACCORD enables complex biological inference using ultrahigh-dimensional multi-omic data through the example of the LIver Hepatocellular Carcinoma (LIHC) cohort ($n$=365) of TCGA. In this analysis, we aim to identify the determinants of mRNA-level expression regulation of individual genes across the tumor samples and compare the strength of different regulatory drivers, particularly between the activity of transcription factors (TF) and the repressive potential of epigenomic regulators using partial correlations as the metric of independent contribution. Since the partial correlation between the mRNA expression level of a gene and the level of each regulatory factor given all other data features deconvolutes the strength of each source of direct regulation, this analysis, if computation is scalable to the entirety of the multi-omic data, will dissect relative contributions of individual regulatory mechanisms to the observed gene expression variation, and the causal graph mirroring the regulatory network is expected to be embedded within the reported (undirected) graph structure.
@@
We used the LIHC dataset consisting of expression data of 15,598 protein-coding genes (mRNAs), expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands positioned in the upstream regulatory regions of genes. These are a subset of variables provided in the original data from the TCGA data portal [1], where only the DNA methylation probes in regulatory regions or gene bodies of annotated genes were included. Since DNA hypo- and hyper-methylation status tends to be correlated locally in genomic neighborhoods and those probesets for methylation status make up a majority of the variables in the data, the underlying graph is expected to reflect a sparse precision matrix with a large number of small block diagonals if the variables were ordered by chromosomes and by within-chromosomal genomic coordinates. Most causally implicated correlations between methylation probes and mRNAs are also likely to be local with the exception of DNA methyltransferase genes and the likes, and they would not alter the presumed network topology described above. The same logic also applies to mRNA-mRNA correlations, but these correlations can be local and global across genomic locations. However, these latter correlations tend to form block diagonal patterns by their upstream TF regulators and the target precision matrix will still remain within a structure amenable to consistent estimation.
@@
We performed this analysis on two HPC systems: Nurion at KSC (see §5.2) and the Cori system at the National Energy Research Scientific Computing Center (NERSC), which is a Cray XC40 system with more than 2,000 computational nodes that are equipped with dual-socket 16-core Intel Xeon Processor E5-2698 (Haswell) and 128GB of memory for each node. On both systems, the shared-memory, multi-core implementation of BigQUIC was the only working alternative for comparison. However, BigQUIC was not able to complete the computation of the precision matrix for a wide range of the regularization parameter $\lambda$, except for those that yielded diagonal matrices; this observation is consistent with the results from the numerical studies in §5.2.
@@
In the next two sections, we first describe the overall partial correlation patterns within and across the two omics data. Using the precision matrix estimate and partial correlations, we aim to tease apart the impact of active TF-driven co-regulation of target genes from that of DNA methylation-mediated gene expression repression in the promoter regions. To show the advantage of this analysis, we compare the mRNA-mRNA network from the two analyses, and show that the analysis accounting for the methylation-mediated regulation recovers experimentally validated regulatory actions of TFs with comparable sensitivity and better specificity. To the best of our knowledge, the effort to deconvolute the effects of two or more types of gene expression regulation using graphical models has not been attempted due to the computational bottleneck we address in this work.
6.2 Network structure of epigenomics and transcriptomics data
@@
We estimated partial correlations by solving (3.2) at $\lambda=0.45$ selected by epBIC (3.7) and then applying the debiasing procedure (3.8). We extracted non-zero partial correlations to form edges and calculated the frequency of intra-modality and inter-modality pairs. Figure 3A shows the frequency of each type of inter- and intra-modality edges, with separate positive and negative edge counts.
Since the DNA methylation probes outnumbered both mRNA- and miRNA-level expression variables in the input data, the majority of the selected edges were the intra-modality correlations between methylation probes ($>$99%), which showed predominantly positive partial correlations (98%) even at long genomic distances. The second most frequent edges were between mRNAs and between miRNAs, where the vast majority showed positive partial correlations (96%), and the correlations were close to zero for the pairs associated with negative partial correlations. This observation clearly reaffirms that physically proximal protein coding genes and non-coding RNAs are often co-transcribed. Among the small number of non-zero inter-modality correlations, the most pronounced category was that of edges between DNA methylation probes and mRNA expression (total transcript level per gene); the signs of the partial correlations were evenly split between positive and negative (Figure 3A).
When we further examined the genomic distances between connected nodes across the six categories, positive inter-modality correlations were from the data features located on the same chromosomes and within 200 kilobase distance of one another (Figure 3B). The edges connecting mRNAs and miRNAs showed exclusively positive partial correlations within a 100 kilobase distance only, hinting at high local specificity of co-transcription. Last but not least, the edges connecting mRNAs and DNA methylation probes showed both positive and negative partial correlations. A closer examination revealed that most large negative partial correlations were between mRNAs and DNA methylation probes within a 1 kbp distance (from TSS), consistent with the established role of methylation for gene expression repression in regulatory regions.
When we further investigated biological functions enriched in the genes with negative correlation between DNA methylation and mRNA expression, those were mostly the genes encoding subunits of enzymes involved in small molecule metabolism and redox reaction regulation (Figure 3C), suggesting that epigenomic regulation actively co-represses the expression levels of metabolic enzymes in liver tumors.
6.3 Dissecting active regulatory contribution of TFs and repressive epigenomic regulation

### v3

@@ 4.2 Edge selection and sign consistency
With further assumptions, the element-wise ($\ell_{\infty}$) error bound of $\hat{\boldsymbol{\Omega}}$ can be controlled, with which edge selection and sign consistency can be proved. The Hessian matrix of the population risk $R(\boldsymbol{\Omega})$ (see §3) with respect to the usual vectorization of $\boldsymbol{\Omega}$ at $\boldsymbol{\Omega}^{*}$ is
$\boldsymbol{\Gamma}^{\ast}=\boldsymbol{\Sigma}^{*}\otimes\boldsymbol{I}_{p}+(\boldsymbol{\Omega}_{D}^{\ast-1}\otimes\boldsymbol{\Omega}_{D}^{\ast-1})\boldsymbol{\Upsilon}.$ (4.1)
where $\otimes$ is a Kronecker product and $\boldsymbol{\Upsilon}=\sum_{i=1}^{p}e_{i}e_{i}^{T}\otimes e_{i}e_{i}^{T}$. Let $S$ be the support of $\boldsymbol{\Theta}^{*}$ as stated in Theorem 4.1. We assume the following property of the data distribution.
Assumption 4.1 (Irrepresentability condition).
There exists $\alpha\in[0,1)$ that satisfies
@@
Let us define the following associated quantities: $\kappa_{\Gamma^{\ast}}=\vvvert(\boldsymbol{\Gamma}^{\ast}_{SS})^{-1}\vvvert_{\infty}$, $\gamma_{1}=\vvvert\boldsymbol{\Omega}^{\ast-1}_{D}\vvvert_{\infty}$. Also, recall that $\kappa_{\Omega^{\ast}}=\vvvert\boldsymbol{\Omega}^{\ast}\vvvert_{\infty}$. These quantities are defined to quantitatively measure and track the model complexity and are allowed to grow along with $(n,p,d)$, where $d$ denotes the maximum number of non-zero entries per row in $\boldsymbol{\Theta}^{*}$.
@@
    there holds $\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{\ast}\|_{\infty}\leq 24\sqrt{2}(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\kappa_{\Gamma^{\ast}}\kappa_{\Omega^{\ast}}(1+10/\alpha)\sqrt{(\tau\log{p}+\log{4})/n}$
    and $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\ast}\|_{\infty}\leq(7/3)\kappa_{\Omega^{*}}\|\hat{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{\ast}\|_{\infty}$, where $\hat{\boldsymbol{\Theta}}=\hat{\boldsymbol{\Omega}}_{D}\hat{\boldsymbol{\Omega}}$ ;
@@
With other quantities held fixed, the sample size required to achieve the rate $\|\hat{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{*}\|_{\infty}=O_{P}(\sqrt{(\tau\log p)/n})$ is $n\gtrsim d^{2}\tau\log p$. It can be also shown that for data distributions with a bounded $4m$-th moment, the error rate is $O_{P}(\sqrt{p^{\tau/m}/n})$ for $n\gtrsim d^{2}p^{\tau/m}$; see Supplementary Material A.6. These sample complexities and rates of convergence match those of graphical lasso, obtained by Ravikumar et al. [33] under similar conditions.
@@
[…]
It follows that the bias-corrected estimator in (3.8) is also consistent:
Corollary 4.1.
Assume the same conditions as Theorem 4.3. Then, for the bias corrected estimator $\breve{\boldsymbol{\Omega}}$ in (3.8), it holds that $\{\operatorname{sign}(\omega_{ij}^{*})=\operatorname{sign}(\breve{\omega}_{ij})\text{ for all }(i,j)\}$ and
$\|\breve{\boldsymbol{\Omega}}-\boldsymbol{\Omega}^{*}\|_{\infty}\leq 24\sqrt{2}(1+4\sigma^{2})(\max_{i\in[p]}\Sigma_{ii}^{*})\kappa_{\Omega^{\ast}}(1+10\alpha^{-1})\sqrt{(\tau\log p+\log 4)/n}$
with a probability no smaller than $1-p^{-(\tau-2)}$
@@ 5.1 Linear convergence of ACCORD-FBS
ACCORD-FBS was investigated with two variants: one with a constant step size of $\tau={1}/{L}$ and the other employing backtracking line search with $\tau={1}/{L}$ as a lower bound. In contrast, a grid of constant step sizes, $\tau\in\{0.25,0.6\}$, along with backtracking, were chosen for ACCORD-ISTA to demonstrate its convergence behavior.
@@ 5.3 Edge detection and estimation error
We examine the edge detection and parameter estimation performance of ACCORD on simulated data, comparing it with CONCORD, graphical lasso (GLASSO), SPACE, CLIME. Following the hub graph setting in Peng et al. [32], we generated a graph with 10 clusters each having 100 nodes and 90 edges and connected 100 randomly chosen node pairs from adjacent clusters. The clusters had either a hub network or scale-free graph structure. A hub network was generated by (1) constructing an Erdös-Rényi graph with 97 nodes and 45 edges, and (2) consecutively adding 3 hub nodes, each connected to random 15 nodes. A scale-free graph was generated so that the degree $k$ of the nodes has a probability $P(k)\sim k^{-2.3}$. We additionally considered a simple Erdös-Rényi graph with 1,000 nodes and 1,000 edges without the clustered structure. The partial correlation matrix corresponding to the graph was chosen so that the matrix has minimum eigenvalue of at least 0.2, and each nonzero entries have absolute value of 0.1 at least. Model selection was carried out using epBIC (3.7) for ACCORD and CONCORD for an evenly spaced grid of $\lambda$’s in logarithmic scale, while extended Gaussian BIC was used for GLASSO and CLIME. Although cross validation was used to select the $\lambda$ in the original paper of CLIME [4], we used BIC instead to avoid the repetitive parameter fitting. For SPACE, we used “BIC-type criterion” defined in Peng et al. [32]. For CONCORD, the matrix $\hat{\boldsymbol{\Theta}}$ minimizing (2.2) was treated as a precision matrix despite the possible inconsistency (see §2). This procedure was repeated 50 times. For ACCORD, we also considered the debiasing procedure (3.8).
@@
Table 1 reports the area under the precision-recall curve (AUPRC) and number of edges selected. In terms of AUPRC (and Matthews correlation coefficient; see Supplementary Material B), ACCORD performed slightly better than or similarly to other methods. The trend of the precision-recall curves did not particularly vary among replications. In all cases, the penalty selected from each selection method yielded reasonable number of selected edges. We could observe that BIC-type methods with the extended term drastically reduces false positive (FP) edges in the expense of few true positive (TP) edges. Table 2 reports the total squared error of the estimated precision matrix $\boldsymbol{\Theta}$ and its reparameterization $\boldsymbol{\Omega}$ by ACCORD and CONCORD, along with their debiased refit. Compared to CONCORD, ACCORD estimates clearly showed better results for both $\hat{\boldsymbol{\Theta}}$ and $\hat{\boldsymbol{\Omega}}$ in terms of the estimation error. The debiased refit also improves the estimation performance. For more details about the impact of the debiased refit, see Supplementary Material B.
@@
(a) Chain graph
(b) Star graph
Figure 3: Maximum error of the ACCORD estimator by sample size $n$
5.4 Estimation error dependency on precision matrix
We further provide numerical results that illustrate how the structures and values of the partial correlation matrix can affect the estimation error of ACCORD. Motivated by experiments in Ravikumar et al. [33], we conducted simulation studies under two types of $\boldsymbol{\Theta}^{*}$: chain and star graphs. In both types, we let $\theta_{ii}=1,$ implying that $\boldsymbol{\Omega}^{*}=\boldsymbol{\Theta}^{*}$. In the chain graph setting, we fix $p=120$ and let $\theta_{ij}=\rho>0$ for $|i-j|=1$ and $\theta_{ij}=0$ for $|i-j|\geq 2.$ Thus, the maximum number of non-zero entries per row $d$ is fixed to $2$. With $\rho$ increasing, the $\kappa_{\Omega^{*}}$ and $\kappa_{\Gamma^{*}}$ also increase and the maximum irrepresentability constant $\alpha$ in 4.1 decreases. Note that for $\rho\geq 0.5$, 4.1 holds for any $\alpha\in[0,1)$. In the star graph setting, $d-1$ nodes among $p=200$ nodes are connected only to a single hub node, and there are no other connections. For all the connected edges, partial correlation entries are set as $2.5/(d-1)$ so that $\kappa_{\Omega^{*}}$ remains invariant with $d$. For sufficiently large $d$, both $\kappa_{\Gamma^{*}}$ and $\alpha$ remain nearly constant. In both settings, an ACCORD estimate is computed with $n$ Gaussian samples drawn from the given $\boldsymbol{\Theta}^{*}$, and the penalty coefficient $\lambda$ is set to be proportional to $\sqrt{1/n}$ as suggested by the theory.
In Figure 3 we provide the maximum estimation error observed with various sample sizes $n$ for the two types of graphs. In the chain graph setting (panel (a)), the rates of the estimation error were similar for $\rho\leq 0.4$, while its magnitude increased with $\rho$. For $\rho=0.5$, the estimator could not recover the chain graph structure, and increasing the sample size did not further improve the estimation error, indicating the need for the irrepresentability condition. In the star graph setting (panel (b)), changing $d$ did not significantly affect the estimation error. This observation is consistent with Theorem 4.2.
@@ 6.1 HP-ACCORD to delineate complex mechanisms in gene expression regulation
In this section, we show that HP-ACCORD enables sophisticated biological inference using ultrahigh-dimensional multi-omic data through the example of the LIver Hepatocellular Carcinoma (LIHC) cohort ($n$=365) of TCGA. We used the LIHC dataset consisting of the expression data of 15,598 protein-coding genes (mRNAs), the expression data of 364 microRNAs (miRNAs) and DNA methylation levels ($\beta$-values) at 269,396 CpG islands located within the upstream regulatory regions of genes. The methylation data contained a subset of variables provided in the original HumanMethylation450k BeadChip array data from the TCGA data portal [1], where we included the probes in genomic regulatory regions or gene bodies only.
In this analysis, the main goal is to identify gene modules whose expression levels are co-regulated by common transcription regulators such as transcription factors (TF) and other co-activators. The graph identified in this analysis can, in turn, facilitate the prioritization of potential expression regulators for downstream biological inference in the context of liver cancer. Epigenetic modulation is an important confounder for this type of analysis. For instance, two genes targeted by a common TF may not show consistent changes in mRNA expression if their epigenetic states were not equally favorable for active transcription. Although DNA methylation does not capture all aspects of the complex epigenetic mechanisms, it is one of the main contributors to this process. Here we aim to estimate partial correlation network from an integrated data set consisting of DNA methylation and mRNA expression. By considering DNA methylation profiles in the same analysis, it is possible to identify gene-to-gene co-regulation network at the mRNA level independent of DNA methylation-driven effects.
@@
Since altered DNA methylation levels tend to be correlated locally in genomic neighborhoods and the probes for DNA methylation represent most variables in the data, the underlying graph is expected to reflect a sparse precision matrix with a large number of small block diagonals when the variables were ordered by genomic coordinates in each chromosome. The most causally implicated correlations between methylation probes and mRNAs are also expected to be from genomic neighborhoods with the exception of DNA methyltransferase genes and others regulating methylation and demethylation processes. By contrast, two genes showing high correlation can be located distantly, or even in different chromosomes. However, these latter correlations often form block diagonal patterns and we expect the target precision matrix to remain within a structure amenable to consistent estimation.
@@
We performed this analysis on two HPC systems: Nurion at KSC (see §5.2) and the Cori system at the National Energy Research Scientific Computing Center (NERSC), which is a Cray XC40 system with more than 2,000 computational nodes that are equipped with dual-socket 16-core Intel Xeon Processor E5-2698 (Haswell) and 128GB of memory for each node. BigQUIC was not able to complete the computation of the precision matrix for a wide range of the regularization parameter $\lambda$, except for those that yielded diagonal matrices; this observation is consistent with the results from the numerical studies in §5.2. As our analysis uses partial correlation as the metric for the relative contribution of epigenetic factors onto gene expression, we decided to forgo CONCORD in this analysis, given the ambiguity in the relationship between its estimate and the precision matrix as discussed in §2.
@@
In the following sections, we describe the partial correlation patterns within and across the two omics data. Using the precision matrix estimate and partial correlations, we aim to tease apart the impact of active TF-driven co-regulation of target genes from that of DNA methylation-mediated gene expression repression in the promoter regions. To the best of our knowledge, the effort to deconvolute the effects of two or more types of gene expression regulation using graphical models has not been attempted due to the computational bottleneck we address in this work. There is a caveat that the effects of other epigenetic regulators such as histone modifications and chromatin states are not accounted for in this analysis. Nonetheless, the analysis clearly shows the advantage of estimation in the ultrahigh-dimensional space when we compare the mRNA-mRNA co-regulation network reported from the two analyses in terms of the percentage of gene pairs with at least one or more shared TFs.
6.2 Network structure of epigenomic and transcriptomic data
@@
We estimated partial correlations by solving (3.2) at $\lambda=0.45$ based on epBIC (3.7) and then applying the debiasing procedure (3.8). We extracted non-zero partial correlations to form edges and calculated the frequency of intra-modality and inter-modality pairs. Figure 4A shows the frequency of each type of inter- and intra-modality edges, with separate positive and negative edge counts.
Since the DNA methylation probes outnumbered both mRNA- and miRNA expression variables in the input data, the majority of the selected edges were intra-modality correlations between methylation probes ($>$99%), which showed predominantly positive partial correlations (98%) at all ranges of genomic distances. The second most frequent edges were between mRNAs and between miRNAs, where the vast majority showed positive partial correlations (96%), and the correlations were close to zero for the pairs associated with negative partial correlations. This observation clearly reaffirms that physically proximal protein coding genes and non-coding RNAs are often co-transcribed [34, 37]. Among the small number of non-zero inter-modality correlations, the most pronounced category was that of edges between DNA methylation probes and mRNA expression (total transcript level per gene); the signs of the partial correlations were evenly split between positive and negative (Figure 4A).
When we examined the genomic distances between connected nodes in the six categories, positive inter-modality correlations were more likely to come from the data feature pairs located on the same chromosomes and within 200 kilobase distance of one another (Figure 4B). The edges connecting mRNAs and miRNAs showed exclusively positive partial correlations within a 100 kilobase distance only, hinting at high local specificity of co-transcription. Last but not least, the edges connecting mRNAs and DNA methylation probes showed both positive and negative partial correlations. A closer examination revealed that the largest negative partial correlations were between mRNAs and DNA methylation probes within a 1 kbp distance (from TSS), consistent with the established role of DNA methylation in the regulatory regions for gene expression repression. We further investigated biological functions enriched in the genes with negative correlation between DNA methylation and mRNA expression. The results showed that the DNA methylation-mediated gene expression repression were mostly observed in the genes encoding subunits of enzymes involved in small molecule metabolism and redox reaction regulation (Figure 4C), suggesting that epigenetic regulation is an active repressor for the gene expression of metabolic enzymes in liver tumors.
6.3 Dissecting active regulatory contribution of TFs and repressive epigenetic regulation